OpenStudy (anonymous):

a sail is in the form of a right triangle that is three times as high as it is wide. the sail is made from 24 square meters of material. what is its height? remember, a=1/4bh please give me the answer like how to right it exactly

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Area = 1/2 base x height 4 square metres = 1/2 base x height but height = 2 x base 4 = 1/2 x base x 2 base 4 = base x base base = square root of 4 = 2 metres Therefore as height = 2 x base, the height = 4 metres

OpenStudy (anonymous):

@iGreen so how exactly would I write that because it is not multiple chice